### v3

# Visual-Neural-Inspired Image Inpainting for Specific Objects-of-Interest Imaging

Abstract
Conventional image inpainting techniques typically process entire images, which often leads to computational inefficiency and susceptibility to information redundancy, particularly in occluded or cluttered scenes. Inspired by cortical processing mechanisms, this study introduces a novel framework termed “Specific Object-of-Interest Imaging” (SIOI) to overcome these limitations. The proposed approach first extracts and encodes object-level representations from complex scenes, producing structural and semantic priors that can be seamlessly integrated into existing inpainting pipelines. Extensive evaluations on dedicated object datasets—Teapot, Elephant, Giraffe, and Zebra—demonstrate that models incorporating SIOI consistently outperform those without it across key metrics including SSIM, PSNR, MAE, and LPIPS. The framework also exhibits strong robustness under challenging conditions such as low illumination, high noise, multiple object occlusions, and motion blur. Furthermore, theoretical analysis grounded in cognitive neuroscience reveals meaningful connections between the “object-first” perceptual mechanism and dynamic feature modulation in visual cortical areas (V1–V4). By significantly enhancing object consistency and semantic coherence across diverse inpainting models, this work not only advances efficient and target-aware image restoration but also fosters interdisciplinary convergence between brain-inspired computation and cutting-edge inpainting technologies.
keywords:
Image Inpainting , Specific Object-of-Interest Imaging , Visual Neural Mechanisms , Deep Neural Network
orrefcempty
[…]
[1]organization=School of Electronic Science and Engineering (School of Microelectronics), South China Normal University, addressline=Foshan, 528000, country=China
[…]
[2]organization=St. Petersburg School of Shipbuilding and Ocean Technology, Guangdong Ocean University, addressline=Zhanjiang 524088, country=China
[…]
[3]organization=Robotics Laboratory, Institute of Automation and Control Processes, addressline=Far Eastern Branch of the Russian Academy of Sciences, city=Vladivostok, postcode=690041, country=Russia
{graphicalabstract}
{highlights}
Research highlight 1
Research highlight 2
1 Introduction
Image inpainting is a fundamental computer vision task that restores missing or corrupted image regions. These imperfections can result from occlusions, sensor noise, data transmission loss, intentional editing, or physical damage. This capability makes it vital for applications like cultural preservation, photography, medical imaging, visual effects, and autonomous systems. However, achieving accurate and semantically coherent inpainting remains challenging.
Conventional inpainting methods primarily rely on intrinsic statistical properties and local structural information of images [1, 2, 3, 4, 5, 6, 7, 8], whose core principle involves leveraging neighborhood information from intact regions to fill missing areas via propagation or copy-paste mechanisms. While adequate for small-scale, texturally regular regions, these methods exhibit significant limitations when confronted with extensive missing areas, complex structural boundaries, or semantically rich regions. They often generate artifacts such as blurred textures, fractured structural continuity, and semantically inconsistent content that undermines global image context.
The rapid development of deep learning—particularly Convolutional Neural Networks (CNNs), Generative Adversarial Networks (GANs), and emerging Transformer- and Diffusion-based models—has driven breakthroughs in complex pattern learning and high-fidelity image synthesis, thereby catalyzing a paradigm shift in image inpainting [9, 10, 11, 12, 13, 14, 15]. These data-driven approaches capture rich image priors and deep semantic representations from large-scale datasets, thereby enabling models to predict missing content with improved accuracy. These models reconstruct intricate textural details while maintaining plausibility.
Despite these advances, significant challenges remain. A key overlooked issue is inefficiency from information redundancy. In practice, inpainting is often object-centric rather than holistic. Consequently, most existing methods process redundant background, incurring substantial overhead. For example, across applications like autonomous driving, reconnaissance, and medical imaging, typically over 90% of data is background rather than target information. Whole-image models consume substantial computational resources analyzing redundant background information, leading to severe resource wastage and reduced accuracy in target-object reconstruction due to background interference, calling for a paradigm shift from holistic processing to targeted, object-focused inference.
In confronting these challenges, biological vision offers an efficient solution. Visual attention mechanisms prioritize relevant targets while suppressing background, enabling efficient target recognition in complex scenes. This ”object-first” approach addresses information sparsity by first localizing targets while excluding background interference. This improves focus, efficiency, and output quality.
Based on this analysis, we propose Specific Object-of-Interest Imaging (SIOI). SIOI decouples object identification from inpainting execution. It serves as a generic front-end that extracts clean object representations while suppressing background, pluggable into any inpainting model. Our framework employs two stages: (1) object perception, and (2) object inpainting. SIOI reduces background interference and focuses processing on target regions, improving efficiency and accuracy. As shown in Tab. 1, this approach aligns with visual cortex hierarchy, mimicking human attentional mechanisms.
Our principal contributions are:
  1. 1.
    We propose Specific Object-of-Interest Imaging (SIOI), a novel framework that decouples object extraction from texture synthesis to reduce background interference and generate structural priors for subsequent tasks.
  2. 2.
    We investigate correlations between human vision and deep networks, integrating neuroscientific insights from visual cortex hierarchy to establish theoretical foundations.
  3. 3.
    We demonstrate SIOI’s effectiveness as a plug-and-play module through comprehensive experiments, showing consistent performance improvements across metrics, datasets, and extreme scenarios.
Table 1: Theoretical Framework Comparison
Model Theoretical Framework Core Mechanism Processing Unit Biological Plausibility
Traditional Classical image processing Handcrafted features Entire image Not considered
Deep-Learning Deep neural networks Learned representations Entire image Not considered
Ours Bio-inspired Object perception → inpainting Specific Object-of-Interest V1-V4 simulation
2 Related Works
Image inpainting has evolved through substantial paradigm shifts, from early mathematical models to modern deep generative architectures. While these approaches differ in implementation, these methods predominantly employ whole-image processing.
2.1 Traditional Mathematical Algorithms: Digital Image Processing and Machine Learning
This period emphasized low-level statistics, local geometric structures, and hand-crafted models. Representative approaches include: 1) PDE- and diffusion-based models [1, 16, 17] that propagate color and gradient information while preserving edges; and 2) patch-based methods [18, 19, 20] that exploit non-local self-similarity by matching and copying patches from intact regions. These approaches [21, 22, 23] reconstruct missing areas through local propagation or patch copying. They perform well on small regular textures but struggle with large occlusions or semantically rich regions due to limited object-level understanding.
2.2 Data-Driven Neural Network Approaches
Deep learning methods improve structural coherence and realism by learning high-level priors. Key paradigms include: (1) CNN encoder-decoder architectures [24, 4, 25, 26, 27] using multi-level features to generate content, with attention mechanisms enhancing focus; (2) GAN-based approaches [28, 29, 30, 31, 32, 33, 34] using adversarial training to produce realistic outputs and fine textures; (3) Transformer-based methods [4, 35, 36, 37, 38, 39, 40] leveraging self- and cross-attention to model long-range dependencies and capture global semantics; and (4) Diffusion models [41, 42, 43] employing progressive denoising to recover image details with fine-grained control and high fidelity.
Yet, most still operate on the entire image, wasting computation on irrelevant background and suffering from interference when target and background overlap. Crucially, they rarely encode explicit object guidance, limiting their ability to ensure object consistency in inpainting. Whole-image models thus suffer from several issues in object-centric tasks: 1) information redundancy and computational inefficiency — processing the entire image wastes computation on irrelevant background content, exacerbating the low effective information density typical in domains such as autonomous driving and satellite imaging; 2) background interference — complex or misleading backgrounds can corrupt the reconstruction of target object structure and semantics, especially when target and background share similar textures or colors; and 3) lack of target-specific guidance and focus — current methods rarely model explicitly which object should be targeted and therefore struggle to separate target signals from background noise during inpainting. Although some methods adopt user sketches or bounding boxes as guidance[44, 45, 46], these require extra interaction and still struggle to achieve precise target separation in complex scenes.
Therefore, a biologically inspired, target-focused imaging paradigm that actively suppresses irrelevant background information is essential for object-centric tasks. Our method, Inpainting via Specific Object-of-Interest Imaging, directly addresses this gap. It introduces a front-end Specific Object-of-Interest Imaging module that algorithmically maximizes target-relevant information and suppresses background interference. This produces refined target-centric signals for subsequent stages, mitigating information redundancy and background interference.
2.3 Insights from Human Visual Mechanisms for Image imaging
[…]
In summary, despite advances in realism, existing methods retain a whole-image bias ill-suited for object-centric tasks. Inspired by biological vision, we propose Specific Object-of-Interest Imaging to address this gap.
3 Proposed Methodology
This section presents our novel image inpainting framework. While conceptually sharing the object-focus goal with other techniques [63, 64, 65, 66, 67], our methodology is fundamentally distinct. The processing pipeline is illustrated in Fig. 1.
Figure 1: Specific Object-of-Interest Imaging technology.
3.1 Specific Object-of-Interest Imaging
Also termed Attention-Targeted Imaging, SIOI builds on neuroscientific foundations. Desimone and Duncan’s model [68] posits visual cortex competition among objects for activation resources. Attention acts as a bias signal amplifying targets while suppressing background, providing a neuroscientific basis for computational background suppression and object enhancement. O’Craven et al. [69] fMRI studies show objects as fundamental attentional units with holistic enhancement. Studies [70, 71] show object boundaries automatically attract attention, engaging V1-V4 regions. Perry et al. [72] reported foveal magnification increases encoding density and task-dependent plasticity enables high-resolution perception. Carrasco and Barbot [73, 74] characterized attention’s dual regulation: exogenous enhances frequency sensitivity, endogenous reallocates resolution. They further showed attentional guidance amplifies high-frequency recognition, enabling region-adaptive reconstruction. Ekman et al. [75] showed attentional signals propagate along V1 object boundaries, inducing feature enhancement—translating to network frequency-domain focusing. Collectively, these reveal object-based prioritization: preferential resource allocation to enhance target representation. These directly inspired our object-detection-guided saliency enhancement applying high-frequency weighting to targets. We integrate saliency modeling with bio-inspired object guidance to construct a cognitively plausible inpainting system (Fig. 2).
Figure 2: Image inpainting framework based on the visual neural Specific Object-of-Interest Imaging.
Anatomically, the visual system processes information hierarchically: local feature detection (retina) $\rightarrow$ spatial integration (LGN) $\rightarrow$ orientation selectivity (V1) $\rightarrow$ holistic object representation (IT cortex). This hierarchical processing suggests that a natural image $k$ comprises multiple latent objects, where V4-mediated attentional modulation enhances target features and suppresses background information. Based on neuronal population-coding and Hebbian plasticity, we model this composition through linear superposition:
$k=\sum_{i=1}^{M}{\{}k_{i}\}$ (1)
where $k\in\mathbb{R}^{H\times W\times C}$ is the scene tensor, $k_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times C}$ represents the $i$-th target object, and $M\in\mathbb{Z}^{+}$ is the total object count. Each $k_{i}$ represents a semantically coherent entity (e.g., a pedestrian or vehicle). Amodal completion focuses on reconstructing fully occluded object structures. The core objective is precise object separation—extracting specific targets $k_{i}$ ($i\in\{1,\dots,M\}$) from cluttered backgrounds. This mirrors the “cocktail-party” effect, where humans focus on a target speaker amidst noise.
To overcome the limitations of traditional methods in noise, we introduce a bio-inspired dynamic feature modulation approach. Its mathematical basis is a target-sensitive measurement matrix $\Gamma$. Through differentiable learning, $\Gamma$ produces a multi-scale feature set $\{\xi_{j}\}_{j=1}^{N}$, capturing the hierarchy from V1 simple cells (edge/texture) to IT complex cells (shape representation). We formalize these concepts next.
For a target index $\mu\in\{1,\dots,M\}$, we have:
$\Gamma k_{\mu}=\sum_{j=1}^{N}\omega_{j,\mu}\xi_{j},$ (2)
where $\Gamma\in\mathbb{R}^{N\times D}$ is the measurement matrix, $\xi_{j}\in\mathbb{R}^{D}$ are the basis vectors, and $\omega_{j,\mu}\in\mathbb{R}$ are the target-specific weights. When optimized, the weights $\omega_{j,\mu}$ concentrate on patterns for $\mu$, creating a selective-gain effect where scene measurements are dominated by the target. This leads to the scene-level measurement:
$\Gamma k=\begin{bmatrix}\sum_{m=1}^{M}\eta_{1,m}\xi_{1}\\ \vdots\\ \sum_{m=1}^{M}\eta_{N,m}\xi_{N}\end{bmatrix}\approx\widehat{k}_{\mu},$ (3)
where $\eta_{j,m}$ is the contribution of object $m$ to channel $j$. Measurements focus on the target when $\eta_{j,m}\to\delta_{m\mu}$ for all $j$, meaning non-target contributions vanish.
Functionally, this models the transformation from V1 (edges/textures) to IT (shapes), with $\Gamma$ acting as an attentional gain in V4. V1 encodes local orientations, while V4 integrates global contours, aligning with our frequency-domain (FFC) focus. From a functional view, this process is a target-guided modulation on a Riemannian manifold:
$\Gamma k=\sum_{m=1}^{M}\Gamma\omega_{m}\otimes k_{m}\approx\hat{k}^{\prime}_{0}$ (4)
where $\otimes$ is the diffeomorphism operator in feature space. A rigorous mathematical model is essential for precise object separation. For a scene with $N$ latent targets, each object admits a sparse representation. Each object $k_{i}$ ($i=0,1,\dots,N-1$) is sparsely represented in a high-dimensional basis $\Psi$:
$k_{i}=\sum_{j=1}^{M}\psi_{ij}x_{ij}$ (5)
where $\psi_{ij}$ is the adaptive basis and $x_{ij}$ the sparse coefficients. The scene decomposes into a linear combination of objects:
$k=\sum_{i=0}^{N-1}k_{i}=\sum_{i=0}^{N-1}\sum_{j=1}^{M}\psi_{ij}x_{ij}$ (6)
subject to sparsity ($\|\mathbf{x}_{i}\|_{0}\ll M$), basis constraints ($\psi_{ij}\in\mathbb{R}^{D}$), and coefficient constraints ($x_{ij}\in\mathbb{R}$). This compressed sensing formulation ensures object separability in feature space.
We implement this via a bio-inspired neural architecture. The Feedforward Feature Selection Network mimics the ventral (”what”) pathway, filtering masked inputs $M_{h}$. The Target Reconstruction Network emulates parietal spatial integration for target reconstruction. Training solves a constrained optimization problem:
$\{\xi,\theta\}=\arg\min\frac{1}{H}\sum_{h=1}^{H}\mathcal{C}(N_{h},\tilde{N}_{h})$ (7)
where $\mathcal{C}(\cdot)$ is the loss function. At convergence, FECNet weights $\xi$ form a feature filter $\phi$. Applying $\phi$ gives:
$\phi k=\phi\left(\sum_{i=0}^{N-1}\sum_{j=1}^{M}\psi_{ij}x_{ij}\right)=\sum_{j=1}^{M}\psi_{0j}x_{0j}=\hat{k}_{0}$ (8)
As a linear operator, $\phi$ has the projection property:
$\phi(x_{ij})=\begin{cases}x_{0j},&\text{if }i=0\\ 0,&\text{if }i\neq 0\end{cases}$ (9)
Assuming the dictionary $\Psi$ satisfies the Restricted Isometry Property (RIP), i.e., $\mathrm{RIP}(\Psi)\leq\delta_{2s}$ (separability condition), the filter $\phi$ approximates an orthogonal complement in feature space ($\phi=I-P_{\mathrm{bg}}$), suppressing non-target components ($i\neq 0$).
A closed-form estimator for $\phi$ is derived by minimizing a least-squares criterion using the Moore-Penrose pseudoinverse:
$\phi=\arg\min_{\phi^{\prime}}\sum_{h}\Big\|\hat{k}_{0}^{(h)}-\phi^{\prime}\Big(\sum_{i=0}^{N-1}k_{i}^{(h)}\Big)\Big\|_{F}^{2}$ (10)
$\Rightarrow\quad\phi=\Big(\sum_{h}\hat{k}_{0}^{(h)}\,k^{(h)\,T}\Big)\Big(\sum_{h}k^{(h)}k^{(h)\,T}\Big)^{\dagger}$
where denotes the Moore-Penrose pseudoinverse, and $k^{(h)}\in\mathbb{R}^{D\times 1}$ is the vectorized scene for sample $h$. We constrain $\phi=\mathbf{I}-\mathbf{P}_{\mathrm{bg}}$ as an approximate orthogonal projection onto the target subspace.
Equivalently, $\phi$ is the best approximation to $I-P_{\mathrm{bg}}$ in a Hilbert space $\mathcal{H}$:
$\phi^{\star}=\arg\min_{\phi\in\mathcal{H}}\;\big\|\,\phi-(I-P_{\mathrm{bg}})\,\big\|_{F}.$ (11)
Here $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and norm $\|\cdot\|$; $\|\cdot\|_{F}$ is the Frobenius norm, where $\|A\|_{F}=\sqrt{\sum_{i}\sum_{j}|a_{ij}|^{2}}$.
Let $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ be learned query and key vectors at locations $(x,y)$ and $(i,j)$, respectively, with $d$ as the feature dimensionality for softmax scaling. Under bio-inspired attentional modulation, attention weights are computed as:
$\alpha_{xy}=\frac{\exp\!\big(\langle\mathbf{q}_{xy},\mathbf{k}_{ij}\rangle/\sqrt{d}\big)}{\sum_{i,j}\exp\!\big(\langle\mathbf{q}_{xy},\mathbf{k}_{ij}\rangle/\sqrt{d}\big)},$ (12)
where the denominator normalizes over all locations $(i,j)$. When $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ are similar, $\alpha_{xy}$ approaches 1, consistent with Feature Integration Theory: biased competition in V4 enhances targets via attentional gain.
We require the Kullback-Leibler (KL) divergence between target and background distributions to satisfy:
$D_{\mathrm{KL}}\big(P_{\mathrm{target}}\|P_{\mathrm{background}}\big)>\sum_{i=2}^{M}D_{\mathrm{KL}}\big(P_{\mathrm{target}}\|P_{\mathrm{object}_{i}}\big),$ (13)
where $D_{\mathrm{KL}}(P\|Q)$ is the KL divergence.
By large-deviation principles, if the target-background discrepancy dominates inter-object interference, empirical distributions concentrate on entropy-minimizing distributions. Thus, the estimator concentrates on the target distribution, reducing estimated target entropy under suitable conditions. Hence, under these assumptions, a unique solution $\widehat{k}_{0}$ exists such that:
$\lim_{t\to\infty}H\big(\widehat{k}_{0}^{(t)}\big)=H_{\mathrm{target}},$ (14)
where $H(X)$ is the Shannon entropy.
Under Assumptions A1–A3 (object sparsity, RIP on $\Psi$, and bounded noise), a linear operator $\phi$ (learnable by FECNet) approximates the ideal projector $I-P_{\mathrm{bg}}$. By compressed-sensing theory, if $\mathrm{RIP}(\Psi)\leq\delta_{2s}$, the learned projector $\phi$ admits the error bound:
$\|\phi k-\widehat{k}_{0}\|_{2}\leq C_{1}\,\sigma_{s}(k)+C_{2}\,\epsilon,$ (15)
where $C_{1}$ and $C_{2}$ depend on RIP constants and the dictionary condition number; $\sigma_{s}(k)$ is the best $s$-term approximation error; and $\epsilon$ bounds the measurement noise.
These results establish a rigorous foundation for our object-centric imaging approach. By formulating target separation as a sparsity-constrained projection problem and using information-theoretic arguments, we show that under our assumptions, the SIOI system provably isolates the target and converges to an optimal reconstruction, up to approximation and noise errors.
3.2 Image Inpainting System
Figure 3: Processing flow of the proposed image-inpainting system.
We propose a two-stage framework based on specific-object-of-interest imaging to enhance structural coherence and reconstruction accuracy for occluded images (Fig. 3). Unlike traditional RGB-space pixel completion, our architecture adopts a biologically inspired object-first approach, decomposing the task into structure-prioritized refinements. We decompose the pipeline into: (1) object perception $\rightarrow$ (2) object inpainting, emphasizing semantic coherence. The core innovation, SIOI, directs model focus to the target while suppressing background interference. This reduces computation by concentrating on critical target regions (see Fig. 4).
Figure 4: Architecture of the proposed Specific Object-of-Interest Imaging.
Stage I: Specific Object-of-Interest Imaging Biological vision starts with retinal phototransduction. Subsequent LGN processing integrates spatial information and suppresses high-frequency noise through center-surround antagonism. In primary visual cortex (V1), simple cells extract oriented edges and spatial-frequency textures [49]. Object prioritization amplifies targets via bottom-up attention [68], suppressing background interference—analogous to the ”cocktail-party” effect.
Our SIOI module computationally implements this hierarchy. Using modified Encoder-Decoder, we process corrupted inputs to extract multi-scale features, mimicking V1 orientation selectivity. FFC modules in encoding/decoding capture long-range dependencies via spectral transformation, simulating V4 attentional modulation that amplifies target signals. The spectral processing is:
$\mathcal{F}_{\mathrm{freq}}=\mathrm{Conv}_{\mathrm{freq}}\big(\mathrm{FFT2}(\mathcal{F})\big),\qquad\mathcal{F}^{\prime}=\mathrm{IFFT2}(\mathcal{F}_{\mathrm{freq}}),$ (16)
where $\mathrm{FFT2}$$\mathrm{IFFT2}$ are 2-D Fourier transform/inverse, and $\mathrm{Conv}_{\mathrm{freq}}$ is a learnable frequency-domain convolutional operator (Fig. 4).
The SIOI module outputs a preliminary but structurally coherent object map with local defects, as shown in Fig. 4. Biologically, this mimics V1 edge detection and V4-mediated attention, where attentional gain enhances targets and suppresses background activity [76]. Our FFC modules amplify target-relevant spectral components and attenuate background spectra. Functionally, FFC operations resemble V1 orientation detection, while attention weights emulate pulvinar-thalamic cortical modulation. Thus, Stage I generates structurally coherent object maps despite occlusions, providing essential priors for texture synthesis and refinement. This spectral processing improves upon standard convolutions by modeling long-range dependencies and enhancing structural consistency.
These mechanisms collectively reduce information redundancy. The SIOI module acts as a front-end visual processor, transforming complex, occluded scenes into purified, object-centric representations. It provides a key intermediate representation, containing structural and semantic priors essential for high-fidelity inpainting.
Stage II: Object Inpainting (Coarse + Fine) To validate SIOI module efficacy independently, we use a simple inpainting network in Stage II. Unlike conventional methods, our model requires no manual masks and automatically identifies and inpaints target corrupted regions. Critically, the innovation is the SIOI module, not this network. Stage II demonstrates that SIOI priors enable high-quality reconstruction even with a simple network. This stage analogizes higher visual areas integrating semantic information for detailed perception. Our Structural Recovery Module (SRM) processes SIOI output with multi-scale dilated convolutions and feature-pyramid fusion for contour interpolation and shape completion:
$\mathcal{F}_{\mathrm{MSDC}}=\mathrm{Concat}\left(\bigcup_{r\in\{1,2,4,6\}}\mathrm{Conv}_{r}(x)\right)$ (17)
FPN-like fusion integrates hierarchical features, and subpixel-convolution upsampling suppresses checkerboard artifacts. The output is a complete structural feature tensor $\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W}$.
Figure 5: Architecture of the Simple Object Inpainting.
Functionally, both systems achieve structural coherence: biological vision uses predictive-coding for shape priors [53], while SRM fuses multi-scale context via MSDC and feature-pyramid fusion. This design avoids texture blurring and semantic mismatches. Technically, MSDC corresponds to V2–V3 contour integration, and feature-pyramid fusion approximates cortical feedback. SRM enhances structural integrity in occluded maps from Stage I, enabling coarse-to-fine imaging.
$\mathcal{F}_{\mathrm{SRM}}=\mathcal{R}(I_{\mathrm{coarse}}),\qquad\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W},$ (18)
where $\mathcal{R}(\cdot)$ is SRM processing and $I_{\mathrm{coarse}}$ is the coarsely restored map. The Global Detail Recovery Module (GDRM) then uses a symmetric encoder-decoder with skip connections to synthesize fine textures, guided by Stage I priors. GDRM implements a symmetric encoder-decoder that consumes the 64-channel features $\mathcal{F}_{\mathrm{SRM}}$ from SRM. This design underscores that SIOI is a versatile, plug-and-play component that elevates any inpainting pipeline via superior object-specific guidance.
3.3 Loss function
To achieve high-fidelity reconstruction of defective object images, we propose a multi-branch composite loss function that simultaneously enhances structural coherence, fine-detail recovery, and global style consistency. It integrates three components: (1) a weighted reconstruction loss [51, 77], (2) a perceptual loss [78], and (3) a style loss [79]. With carefully designed weighting schemes and a multi-stage optimization strategy, this composite loss ensures semantically coherent structures and texturally faithful details in the synthesized images.
Weighted Reconstruction Loss. We first introduce a spatially weighted reconstruction loss guided by target-specific maps. An explicit weight map $w(x,y)$ differentially weights per-pixel reconstruction errors in the $L_{1}$$L_{2}$ terms:
$\mathcal{L}_{\mathrm{weighted}}\;=\;\frac{1}{N}\sum_{x,y}\big(\alpha\,w(x,y)+1\big)\,\big\|\hat{I}(x,y)-I_{\mathrm{gt}}(x,y)\big\|_{2}^{2}$ (19)
Equivalently, at pixel-level the reconstruction loss can be written as
$\mathcal{L}_{\mathrm{recon}}\;=\;\sum_{i}w_{i}\,\big\|I_{i}^{\mathrm{pred}}-I_{i}^{\mathrm{gt}}\big\|_{2}^{2},$ (20)
where $\hat{I}(x,y)$ denotes the predicted intensity at pixel $(x,y)$, $I_{\mathrm{gt}}(x,y)$ the corresponding ground truth, $I_{i}^{\mathrm{pred}}$ and $I_{i}^{\mathrm{gt}}$ the RGB vectors at pixel $i$, $w_{i}$ the pixel-level weight, and $\alpha$ a scalar weighting factor (empirically set to 4). This weighting scheme prioritizes detail reconstruction in semantically important regions while reducing emphasis on background areas, effectively mitigating edge blurring and structural discontinuities. This approach enhances supervision on target regions while minimizing interference from irrelevant background gradients.
Perceptual Loss. To address the limitation of pixel-level losses in capturing high-level semantics, we incorporate a perceptual loss derived from feature activations of a pretrained VGG-16 network:
$\mathcal{L}_{\mathrm{perc}}\;=\;\sum_{l\in\mathcal{L}}\frac{1}{C_{l}H_{l}W_{l}}\,\big\|\phi_{l}(I^{\mathrm{pred}})-\phi_{l}(I^{\mathrm{gt}})\big\|_{2}^{2},$ (21)
where $\phi_{l}(\cdot)$ denotes the feature maps extracted from the $l$-th VGG-16 layer, and $C_{l},H_{l},W_{l}$ denote the channel count, height, and width of $\phi_{l}$, respectively. By comparing deep feature representations across selected VGG-16 layers, the perceptual loss captures high-level semantic information (e.g., object shapes and textures) and improves semantic coherence in the generated outputs.
Style Loss. The style loss measures texture and style characteristics via Gram matrices, following style-transfer approaches [79]. For feature maps $\phi_{l}(x)$ at layer $l$ of VGG-16, the Gram matrix is
$G_{l}(x)\;=\;\phi_{l}(x)\,\phi_{l}(x)^{\top},$ (22)
and the style loss is
$\mathcal{L}_{\mathrm{style}}\;=\;\sum_{l\in\mathcal{L}}\big\|G_{l}(I^{\mathrm{pred}})-G_{l}(I^{\mathrm{gt}})\big\|_{F}^{2},$ (23)
where $\mathcal{L}$ denotes the set of selected layers and $\|\cdot\|_{F}$ the Frobenius norm. This statistical matching captures both local texture patterns and global style consistency, enhancing realism in restored images while reducing artifacts and repetitive patterns.
Composite Loss. Finally, we combine these components into a composite loss:
$\mathcal{L}_{\mathrm{total}}=\lambda_{1}\mathcal{L}_{\mathrm{recon}}+\lambda_{2}\mathcal{L}_{\mathrm{perc}}+\lambda_{3}\mathcal{L}_{\mathrm{style}}\quad(\lambda_{1}=1.0,\,\lambda_{2}=0.1,\,\lambda_{3}=250)$ (24)
The reconstruction term sharpens edges in key regions, the perceptual term ensures semantic alignment with human vision, and the style term reduces texture artifacts across varied content. These weights balance the varying scales and objectives of the loss components during training.
4 Experiment
4.1 Dataset
We utilized four distinct datasets for our experiments. This included a custom, object-centric teapot dataset (2,000 training and 200 test samples), created in-house. It contains $512\times 512$‘ images of teapots from multiple viewpoints, set against complex backgrounds with artificial items and indoor settings. Each image has a corresponding $512\times 512$‘ ground-truth mask, meticulously calibrated through a multi-stage manual process to ensure accurate geometric and textural details. Additionally, three subsets were derived from COCO2017: “elephant” (2,143 training, 89 validation-test), “giraffe” (2,546 training, 101 validation-test), and “zebra” (1,916 training, 85 validation-test). All images were resized to $512\times 512$‘ to maintain consistency. A consistent occlusion protocol was applied across all datasets. Masks were automatically generated from target maps to ensure occlusions affected only the foreground objects, covering approximately $\sim$40%‘ of their area. This balance preserved sufficient context while requiring the model to perform non-trivial structural reasoning during both training and inference.
4.2 Experimental Setup
Implementation Details. All experiments were run on a Ubuntu 20.04 system with an Intel Xeon Platinum 8358P CPU and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM). We implemented our framework in PyTorch 1.10.0 using CUDA 11.3. Training lasted for 50 epochs with a batch size of 4, a constraint imposed by our computational resources. We used the Adam optimizer with a learning rate of $2\times 10^{-4}$‘. The total training time was approximately 0.5 hours on this hardware. Our training followed a two-stage protocol. First, we pre-trained the ObjectImagingModule using only the reconstruction loss $\mathcal{L}_{\mathrm{recon}}$‘ to produce structured intermediate representations. This plug-and-play module was then frozen and integrated into all subsequent experimental setups. Second, we trained the Structural Refinement Module (SRM) and Geometric Detail Reconstruction Module (GDRM) with the perceptual loss $\mathcal{L}_{\mathrm{perc}}$‘ and style loss $\mathcal{L}_{\mathrm{style}}$‘ to improve semantic coherence and stylistic fidelity. During this second stage, the ObjectImagingModule’s parameters remained fixed. This optimization was conducted primarily for preliminary validation in Experiment 1. Our full implementation code is available at: https://github.com/WYH302/WuYonghao.git.
Baseline. To rigorously evaluate our method, we benchmark against a representative selection of state-of-the-art inpainting approaches. Our selection emphasizes methodological diversity, encompassing classical exemplar-based methods to modern generative approaches, including diffusion models, transformers, and GAN architectures. We include Criminisi [18] as a foundational exemplar-based method that uses non-local patch synthesis, providing a baseline for traditional approaches while revealing limitations in complex structural recovery. PatchMatch [80], an optimized variant, offers insights into computational efficiency through accelerated nearest-neighbor matching. From modern deep learning, AOT-GAN [81] is included for its attention-driven modeling of long-range dependencies using dilated convolutions, addressing holistic scene coherence. Edge-Connect [44] provides a two-stage edge-guided approach that allows comparison of structural prior usage with our coarse-to-fine framework. The hybrid MISF [82] tests multi-scale feature interactions between traditional and deep components, providing a contrast to our bio-inspired approach. Finally, we include contemporary diffusion models: LaMa [83] validates our spectral processing choices, while RePaint [41] provides a stochastic refinement contrast to our deterministic approach. All baseline methods were executed for five independent runs with random initializations, using official implementations and pre-trained models when available.
Evaluation metrics. We used four standard metrics: (1) SSIM measures structural similarity (higher values preferred), (2) PSNR quantifies pixel-level fidelity (higher values better), (3) MAE calculates average pixel deviation (lower values better), and (4) LPIPS [84] evaluates perceptual quality (lower values better).
4.3 Preliminary theoretical verification
To validate our core contribution—the Specific Object-of-Interest Imaging (SIOI) module—we conducted a controlled experiment using a simplified inpainting network (Stage II: SRM+GDRM). This network is intentionally simplified and not our innovation focus. Its purpose is to demonstrate that high-quality outputs primarily stem from SIOI’s structural and semantic priors (Stage I).
Unlike conventional methods, our model automatically identifies and inpaints corrupted regions without manual mask annotation. We retained black borders to show the model doesn’t rely solely on color cues for corruption detection. As shown in Table 2 and Fig. 6, this pipeline achieves exceptional performance on the Teapot dataset. These results confirm the SIOI module successfully extracts clean, structurally sound object representations from complex, occluded scenes. This validates SIOI’s efficacy as a standalone component for downstream inpainting tasks.
Table 2: Preliminary Theoretical Verification on the teapot dataset (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better.
[…]
Experimental values $0.98\pm 0.02$ $33.86\pm 0.38$ $1.61\pm 0.18$ $0.02\pm 0.01$
Figure 6: Preliminary Theoretical Verification of exemplar inpainting results (teapot dataset).
4.4 Generalized empirical evaluation——Quantitative comparisons
We next evaluate SIOI’s utility as a universal plug-and-play front-end for enhancing diverse inpainting models. We integrated SIOI with state-of-the-art methods spanning classical to modern deep learning approaches. For fair comparison, all methods were evaluated using a consistent object-level protocol, with outputs multiplied by corresponding object masks.
As shown in Table 3, SIOI consistently improves performance across all metrics. The key finding is the transformative gain from prepending SIOI, not any single model’s absolute performance. For example, SIOI improves Criminisi’s PSNR by 7 dB to 31.552 dB. Notably, SIOI boosts LaMa to 45.222 dB PSNR and 0.0023 LPIPS, surpassing all other methods. Improvements are most pronounced on the Teapot dataset, with SIOI+LaMa achieving near-perfect reconstruction (SSIM: 0.9973, LPIPS: 0.0023). These results show SIOI enables both traditional and modern algorithms to transcend their limitations through purified, object-centric inputs. SIOI also improves computational efficiency by bridging holistic image processing and efficient object-centric reconstruction.
Across datasets, models with SIOI consistently outperform those without it across all metrics. Models with SIOI also exhibit enhanced perceptual quality. These results confirm SIOI’s effectiveness for high-fidelity reconstruction of complex-textured objects. The object-centric paradigm enables consistent target separation across diverse backgrounds. Computationally, SIOI reduces redundant processing by focusing on target regions, accelerating inference while maintaining quality.
Table 3: Experimental Results for Different Datasets and Models, ↑ means higher is better, and ↓ means lower is better.
Datasets Model SSIM $\uparrow$ PSNR $\uparrow$ MAE $\downarrow$ LPIPS $\downarrow$
Teapot Criminisi 0.9415 24.597 3.5517 0.0965
PatchMatch 0.9212 21.219 5.4619 0.1263
Edge-Connect 0.9457 25.095 3.3687 0.0908
AOT-GAN 0.9487 25.329 3.2564 0.0863
MISF 0.9370 24.871 3.7130 0.1178
LaMa 0.9524 25.534 3.1024 0.0833
RePaint 0.9330 24.037 4.1087 0.1141
SIOI+Criminisi 0.9841(↑4.5%) 31.5520(↑28.3%) 0.7396(↓79.2%) 0.0246(↓74.5%)
SIOI+PatchMatch 0.9752(↑5.9%) 25.9120(↑22.1%) 1.7894(↓67.2%) 0.0408(↓67.7%)
SIOI+Edge-Connect 0.9795(↑3.6%) 32.3780(↑29.0%) 0.8364(↓75.2%) 0.0299(↓67.1%)
SIOI+AOT-GAN 0.9248(↓2.5%) 36.9400(↑45.8%) 1.0471(↓67.8%) 0.0094(↓89.1%)
SIOI+MISF 0.9739(↑3.9%) 33.4870(↑34.6%) 1.0541(↓71.6%) 0.0674(↓42.8%)
SIOI+LaMa 0.9973(↑4.7%) 45.2220(↑77.1%) 0.1119(↓96.4%) 0.0023(↓97.2%)
SIOI+RePaint 0.9712(↑4.1%) 28.5990(↑19.0%) 1.6907(↓58.9%) 0.0645(↓43.5%)
elephant Criminisi 0.9671 32.615 1.3453 0.0262
PatchMatch 0.9596 28.643 2.1005 0.0441
Edge-Connect 0.9562 31.958 1.7455 0.0423
AOT-GAN 0.8416 27.653 4.9353 0.0608
MISF 0.8472 23.789 5.2913 0.1908
LaMa 0.9236 26.059 2.9596 0.0949
RePaint 0.8624 23.690 5.1769 0.1676
SIOI+Criminisi 0.9774(↑1.1%) 33.1520(↑1.6%) 0.8312(↓38.2%) 0.0190(↓27.5%)
SIOI+PatchMatch 0.9721(↑1.3%) 29.3270(↑2.4%) 1.3736(↓34.6%) 0.0325(↓26.3%)
SIOI+Edge-Connect 0.9636(↑0.8%) 30.6930(↓4.0%) 1.2479(↓28.5%) 0.0448(↑5.9%)
SIOI+AOT-GAN 0.9341(↑11.0%) 34.6910(↑25.5%) 1.4350(↓70.9%) 0.0243(↓60.0%)
SIOI+MISF 0.8883(↑4.9%) 29.0700(↑22.2%) 3.4237(↓35.3%) 0.1848(↓3.1%)
SIOI+LaMa 0.9786(↑6.0%) 34.7060(↑33.2%) 0.6523(↓78.0%) 0.0140(↓85.2%)
SIOI+RePaint 0.9102(↑5.5%) 28.9560(↑22.2%) 3.2348(↓37.5%) 0.1489(↓11.2%)
giraffe Criminisi 0.9721 29.962 1.5050 0.0184
PatchMatch 0.9653 27.671 2.3031 0.0368
Edge-Connect 0.9678 29.914 1.6754 0.0280
AOT-GAN 0.9109 26.742 4.1489 0.0388
MISF 0.8998 23.505 4.6733 0.1208
LaMa 0.9424 25.732 2.5901 0.0670
RePaint 0.9079 23.563 4.5075 0.1096
SIOI+Criminisi 0.9788(↑0.7%) 29.8240(↓0.5%) 1.1000(↓26.9%) 0.0170(↓7.6%)
SIOI+PatchMatch 0.9734(↑0.8%) 26.3270(↓4.9%) 1.7493(↓24.0%) 0.0306(↓16.8%)
SIOI+Edge-Connect 0.9676(↑0.0%) 27.5270(↓8.0%) 1.6295(↓2.7%) 0.0382(↑36.4%)
SIOI+AOT-GAN 0.8741(↓4.0%) 30.8610(↑15.4%) 2.3573(↓43.2%) 0.0350(↓9.8%)
SIOI+MISF 0.9325(↑3.6%) 27.2700(↑16.0%) 3.0461(↓34.8%) 0.1215(↑0.6%)
SIOI+LaMa 0.9796(↑3.9%) 31.0230(↑20.6%) 0.9491(↓63.4%) 0.0129(↓80.7%)
SIOI+RePaint 0.9465(↑4.3%) 27.1090(↑15.0%) 2.8830(↓36.0%) 0.0992(↓9.5%)
zebra Criminisi 0.9677 27.598 2.2333 0.0236
PatchMatch 0.9262 22.627 4.1497 0.0857
Edge-Connect 0.9588 27.891 3.1040 0.0263
AOT-GAN 0.8451 22.399 9.6206 0.0486
MISF 0.8583 21.409 8.2618 0.1596
LaMa 0.9331 24.284 3.7468 0.0684
RePaint 0.8677 21.544 8.0205 0.1499
SIOI+Criminisi 0.9750(↑0.8%) 27.3890(↓0.8%) 1.6940(↓24.1%) 0.0211(↓10.6%)
SIOI+PatchMatch 0.9675(↑4.5%) 24.6910(↑9.1%) 2.4553(↓40.8%) 0.0365(↓57.4%)
SIOI+Edge-Connect 0.9632(↑0.5%) 26.0890(↓6.5%) 2.2360(↓28.0%) 0.0387(↑47.1%)
SIOI+AOT-GAN 0.9070(↑7.3%) 28.0250(↑25.1%) 3.2725(↓66.0%) 0.0527(↑8.4%)
SIOI+MISF 0.8960(↑4.4%) 24.0480(↑12.3%) 6.2437(↓24.4%) 0.1556(↓2.5%)
SIOI+LaMa 0.9772(↑4.7%) 28.8110(↑18.6%) 1.3896(↓62.9%) 0.0144(↓78.9%)
SIOI+RePaint 0.9130(↑5.2%) 24.3090(↑12.8%) 5.9439(↓25.9%) 0.1365(↓8.9%)
4.5 Generalized empirical evaluation——Qualitative comparisons
Qualitative results confirm SIOI’s effectiveness as a preprocessing stage. SIOI provides clean, background-suppressed object representations, enabling conventional algorithms to produce more consistent and plausible results. As shown in Figures 7-14, benchmark models exhibit characteristic failure modes: Criminisi [18] lacks semantic understanding, causing structural discontinuities and inconsistent textures; PatchMatch [80] relies on hand-crafted features, producing visible artifacts due to localized patch processing; AOT-GAN [81] shows jagged artifacts from dilated convolutions; Edge-Connect [44] suffers spatial misalignment from over-reliance on Canny edges; MISF [82] produces oversmoothed edges and illumination inconsistencies; LaMa [83] oversmooths details, producing blurry outputs; RePaint [41] yields blurry outputs with incongruous textures. Without SIOI, methods fail due to background interference and lack of object understanding.
With SIOI, all methods show improved edge sharpness, texture consistency, and semantic coherence. Our approach excels on rigid objects (e.g., teapots) but shows relatively weaker results on non-rigid objects (giraffes, zebras, elephants). This difference stems from higher morphological variability in non-rigid categories, requiring larger or more adaptive models. This suggests a relationship between semantic complexity, data scale, and model size, warranting future study.
Figure 7: Qualitative comparison of exemplar inpainting results (teapot dataset).
Figure 8: Qualitative comparison of exemplar SIOI inpainting results (teapot dataset).
Figure 9: Qualitative comparison of exemplar inpainting results (elephant dataset).
Figure 10: Qualitative comparison of exemplar SIOI inpainting results (elephant dataset).
Figure 11: Qualitative comparison of exemplar inpainting results (giraffe dataset).
Figure 12: Qualitative comparison of exemplar SIOI inpainting results (giraffe dataset).
Figure 13: Qualitative comparison of exemplar inpainting results (zebra dataset).
Figure 14: Qualitative comparison of exemplar SIOI inpainting results (zebra dataset).
4.6 Ablation Studies
In fact, a detailed ablation experiment has already been conducted above, So this is just a simple supplement to the Preliminary theoretical verification experiment, also conducted on the Teapot dataset. We conducted additional ablation studies on the Teapot dataset under identical conditions. Key components were systematically removed to analyze performance differences, as shown in Fig. 15.
Figure 15: Illustration of ablation experiment.
Ablation I: Removing Stage I while retaining object-centric I/O (Tab. 4) showed marginal metric improvements (SSIM: 0.976 vs. 0.978, PSNR: 33.95 dB vs. 33.86 dB, MAE: 1.313 vs. 1.605, LPIPS: 0.017 vs. 0.018), but these were statistically insignificant. Subjectively, edge localization degraded with jagged artifacts and blurring, confirming Stage I’s critical role for structural precision. The metric improvements, despite perceptual degradation, highlight that pixel-based metrics can mislead due to their focus on RGB deviations rather than human perception.
Ablation II: Replacing object-centric I/O with full-scene inputs (standard paradigm) caused severe performance degradation: SSIM: 0.964, PSNR: 29.43 dB, MAE: 2.085, LPIPS: 0.065. Visual inspection showed poorest quality with detail loss, artifacts, and impaired clarity. The model’s inability to extract features from degraded regions led to blurry reconstruction and artifacts. This highlights our object-first advantage: focusing computation on the target avoids whole-image pitfalls.
In contrast, the full model (with SIOI) achieved superior performance with sharp edges, semantic coherence, and photorealistic textures. This confirms that our two-stage pipeline effectively addresses traditional inpainting bottlenecks by first resolving target object localization and identity before inpainting. SIOI serves as a powerful pre-processor that enhances any inpainting network for object-centric tasks.
Table 4: Comparison of ablation study results on the Teapot dataset. (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better (teapot dataset).
[…]
Origin $0.98\pm 0.01$ $33.86\pm 0.38$ $1.61\pm 0.08$ $0.02\pm 0.01$
Ablation I $0.98\pm 0.01$ $33.95\pm 0.19$ $1.31\pm 0.06$ $0.02\pm 0.01$
Ablation II $0.96\pm 0.01$ $29.43\pm 0.25$ $2.09\pm 0.15$ $0.07\pm 0.02$
4.7 Robustness Analysis
This study extends the preliminary theoretical verification experiment on the Teapot dataset. To assess generalization under challenging conditions, we tested robustness in scenarios:
  • 1.
    Low illumination: 50% reduced scene brightness
  • 2.
    Small targets: Objects occupying $5\%$ image area
  • 3.
    Gaussian noise: Additive noise ($\mu=0,\sigma=1$)
  • 4.
    Salt-and-pepper noise: Impulse noise ($p_{\text{salt}}=0.01,p_{\text{pepper}}=0.01$)
  • 5.
    Multi-object occlusion: 3 instances
  • 6.
    Motion blur: Linear kernel with $15\times 15$ pixels
Note that these scenarios were absent from training data. For qualitative assessment, we selected two samples per category. Quantitative evaluation was omitted due to limited samples per condition, as shown in Fig. 16.
Qualitative results show SIOI’s robustness in several challenging conditions. It maintained satisfactory performance under low illumination, small targets, salt-and-pepper noise, and motion blur. This resilience stems from its design: spectral processing (FFC) resists global illumination changes and motion artifacts, while attentional mechanisms, inspired by V4, amplify and track small targets despite noise. Performance degraded with three or more objects due to attentional resource competition. This observation directly mirrors the biological phenomenon of attentional limitation described in Desimone and Duncan’s biased competition model [68].
Severe degradation occurred under high-power Gaussian noise. This challenge is known for feature-based attentional systems in biological and artificial networks. SIOI’s reliance on gradient and frequency information makes it susceptible to non-structured, high-entropy noise, overwhelming salient features. This highlights SIOI’s critical role and shortcoming: its output quality underpins all subsequent processing. Our modular design should contains failures at SIOI stage, preventing system collapse and pinpointing improvements. These findings guide SIOI’s evolution. Future work will enhance robustness against unstructured noise using neuromorphic adaptations from auditory brainstem or retinal gain control. This will solidify SIOI’s role as a plug-and-play pre-processor for object-centric vision.
In summary, SIOI provides reliable object-specific priors across challenging conditions. Its profile—strong against structured degradation but susceptible to unstructured noise—defines its operational envelope and guides application as a bio-inspired front end.
Figure 16: Illustration of robustness experiment (teapot dataset).
5 Discussion
Image inpainting lacks a definitive standard solution and frequently yields contentious results, as illustrated in Fig. 17. It is often ambiguous whether a missing region should be inpainted as one object or another (e.g., a buffalo or zebra head). Furthermore, the same model can produce varying outputs for an identical input across different runs. Our proposed framework addresses this challenge by introducing the Specific Object-of-Interest Imaging (SIOI) paradigm—a bio-inspired approach that reconceptualizes the processing pipeline for object-centric vision tasks. Departing from conventional holistic processing, SIOI provides a plug-and-play front-end module that first isolates and enhances the target object, generating a refined structural and semantic prior for any subsequent processing network. Our experiments demonstrate that incorporating SIOI consistently elevates the performance of diverse inpainting architectures. Thus, our contribution is not to propose a superior inpainting network, but to demonstrate that providing a high-quality, object-centric prior is a more impactful strategy than developing another holistic model.
Figure 17: Illustration of challenges in image inpainting.
This approach mirrors the ventral visual pathway’s ”object-before-background” processing strategy, where attentional mechanisms (e.g., in area V4) enhance relevant features while suppressing background clutter [85]. Our SIOI module computationally implements this principle, mitigating the core issue of informational redundancy inherent in whole-image processing. Unlike GAN- or diffusion-based methods that process the entire scene, our approach avoids error propagation from background regions by concentrating computation on the target object, analogous to the role of V4 in contour integration and feature binding.
The robustness of SIOI under challenging conditions (e.g., low illumination, noise) further highlights its utility as a reliable pre-processing module. Its behavior is consistent with the object-based attention mechanism in biological vision, which dynamically modulates early visual areas (V1-V4) to maintain target salience amidst distraction. Notably, the performance of any downstream model is contingent on the quality of the prior provided by SIOI. This dependency parallels the dorsal-ventral synergy in the brain, wherein spatial localization guides object recognition. This highlights that advancing object-level pre-processing is a critical frontier in itself.
We emphasize that the inpainting network in Stage II serves primarily as a proof-of-concept to validate the efficacy of the SIOI prior. Consequently, our architecture does not represent the performance ceiling, and future work may develop superior networks. The paramount innovation of this work is not this specific network, but the SIOI paradigm itself. This paradigm shift—from direct holistic inpainting to object-first prior generation—is our central contribution. The generalization capability of the SIOI module is inherently linked to the semantic definition of a ”specific object.” This is not merely a limitation but a defining characteristic and a research opportunity. We acknowledge that the current generalization capability is limited due to its focus on specific object categories. Each semantically distinct object category effectively constitutes a separate dataset, necessitating retraining of the object imaging network. The granularity of this definition—for instance, whether ”cat” constitutes a single category or should be divided into subspecies (e.g., Felis chaus, Felis margarita)—requires further investigation. However, the module’s performance can be scaled and specialized, providing a practical pathway for technology transfer and enabling precise tailoring to application-specific needs.
Future work will explore integrating prefrontal-cortex-like, task-driven modulation to enable dynamic reconfiguration of SIOI for different object categories and goals, advancing toward adaptive and biologically plausible vision systems.
6 Conclusion
In conclusion, we propose Specific Object-of-Interest Imaging (SIOI), a bio-inspired preprocessing paradigm for object-centric visual tasks. It operates by isolating and enhancing the target object, providing a high-fidelity prior that enhances subsequent vision modules. Extensive experiments confirm SIOI as a powerful, plug-and-play front-end that improves diverse inpainting models by suppressing background interference and providing a coherent object-centric prior. Grounded in the ventral pathway’s ”object-first” hierarchy, it offers a computational solution to information redundancy. The framework demonstrates robustness under challenging conditions, reflecting the resilience of biological attention. A limitation is its reliance on initial object generation, which can impair boundary guidance. The SIOI paradigm can benefit other object-centric tasks like segmentation, detection, and tracking, where suppressing background clutter is critical. Future work includes task-driven modulation and fMRI validation, advancing toward biologically adaptive imaging for neuroimaging and perceptual assistance.
Table 5: Extended Mathematical Notation Guide with Biological Correlations
Symbol Mathematical Definition Biological Mechanism Network Implementation
$\Gamma$ Target-sensitive measurement matrix: $\Gamma k\approx\sum_{m=1}^{M}(\Gamma\omega_{m})\otimes k_{m}$ LGN attentional filtering SIOI spectral filtering
$k$ Scene tensor: $k\in\mathbb{R}^{H\times W\times C}$ Retinal photoreceptor output encoding SIOI degraded input (Stage I)
$\otimes$ Diffeomorphic operator: $(\Gamma\omega_{m})\otimes k_{m}$ Foveal cortical magnification FFCModule feature warping
$\xi_{j}$ Orthogonal feature basis: $\xi_{j}\in\mathbb{R}^{D}$ V1 simple cell orientation responses Multi-scale encoder features in SIOI
$\omega_{m}$ Feature weights: $\omega_{m}\rightarrow\delta_{m\mu}$ Neural resource allocation Attention weights
$k_{i}$ Latent object tensor: $k_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times C}$ IT cortex semantic object representation Target object map (Stage I output)
$\phi$ Projection operator: $\phi=I-P_{\mathrm{bg}}$ Dorsal pathway spatial suppression TRNet background suppression
$D_{KL}$ Kullback–Leibler divergence: $D_{KL}(P\|Q)=\sum_{i}P_{i}\log\frac{P_{i}}{Q_{i}}$ Neural coding efficiency optimization Loss function convergence
$\mathcal{F}$ Spectral feature map: $\mathcal{F}^{\prime}=\mathrm{IFFT2}\!\left(\mathrm{Conv_{freq}}\!\left(\mathrm{FFT2}(\mathcal{F})\right)\right)$ V1–V4 spatial frequency processing FFCModule operations in bottleneck
$\mathcal{L}_{\mathrm{total}}$ Composite loss: $\mathcal{L}_{\mathrm{total}}=\lambda_{1}\mathcal{L}_{\mathrm{recon}}+\lambda_{2}\mathcal{L}_{\mathrm{perc}}+\lambda_{3}\mathcal{L}_{\mathrm{style}}$ Prefrontal cortex task-driven modulation Phased training strategy